OpenStudy (anonymous):

Solve for x^2 +2x-24=0

OpenStudy (hoblos):

x^2 +2x-24=0 (x+8)(x-6)=0 x=-8 or x=6

OpenStudy (king):

u can also use quadratic equation formulaso, x=-8,6

OpenStudy (anonymous):

what is quad formula again?